Activity-Based Budget
Macon Industries has the following budget information available for February:

Units manufactured 35,000
Factory administration $50,000
Assembly 0.5 hour per unit X $9
Direct materials 2 pounds per unit X $2
Inspection $100 per batch of 1,000 units
Manufacturing overhead $4 per unit
Product development $20,000
Setup cost $11 per batch of 1,000 units


Use activity based costing to prepare a manufacturing cost budget for February. Clearly distinguish between unit, batch, and facility-level costs.

Macon Industries
Activity-Based Manufacturing Cost Budget
For the Month of February
Unit-level costs:
Direct materials $Answer
Assembly Answer
AnswerFactory administrationInspectionManufacturing overheadProduct development Answer Answer
Batch-level costs:
Setup Answer
AnswerFactory administrationInspectionManufacturing overheadProduct development Answer Answer
Product-level costs:
AnswerFactory administrationInspectionManufacturing overheadProduct development Answer
Facility-level costs:
AnswerFactory administrationInspectionManufacturing overheadProduct development Answer
Budgeted manufacturing costs $Answer
Answer #1

Solution

Macon Industries

Manufacturing Cost Budget for February using activity based costing:

Macon Industries

Activity-Based Manufacturing Cost Budget

For the Month of February

Unit-Level Costs:

Direct Materials

$140,000

Assembly

$157,500

Manufacturing Overhead

$140,000

$437,500

Batch-Level Costs:

Setup

$385

Inspection

$3,500

$3,885

Product-Level Costs:

Product Development

$20,000

Facility-Level Costs:

Factory Administration

$50,000

Budgeted Manufacturing Costs

$511,385

Computations:

  1. Direct materials cost –

= 35,000 units x 2 pounds/unit x $2 = $140,000

  1. Assembly –

= 35,000 x 0.5 hours/unit x $9 = $157,500

  1. Manufacturing overhead –

= 35,000 units x $4 per unit = $140,000

  1. Setup cost –

= (35,000/1,000) = 35 batches per x $11 per batch = $385

  1. Inspection costs –

= (35,000/1,000) x $100 = $3,500